Find the optimum level of output and profit from the cost function

TC = 50 + 6Q2

 and price

P = 100 – 4Q

Also derive marginal cost and marginal revenue


1
Expert's answer
2021-04-30T10:51:16-0400

"TC=50+6Q^2"

"P=100-4Q"


"MC=\frac {d(TC)}{dQ}=0+12Q=12Q"


At profit max point; "100-4Q=12Q"


"=100=12Q+4Q"

"=\frac{100}{16}=Q"

Optimal level of output ; "Q=6.25"


ii. Total Revenue =Price"\times"Quantity

At "Q=6.25"


"TR=[100(4\times6.25)]\times6.25"

"=468.75"

Total cost (TC) "=50+6(6.25)^2"

"=284.375"


Therefore; "profit=TR-TC"

"=468.75-284.375"


"Profit=184.375"


Deriving Marginal cost and Marginal Revenue.

Given that; "P=100-4Q"

Multiply both sides by "Q"

"PQ=TR=100Q-4Q^2"

"=MR=\frac{d(TR)}{dQ}=(100\times1)-(4\times2)Q^2-^1"


"MR=100-8Q"



"MC=" "\frac {d(TC)}{dQ}=0+12Q=12Q"


"MC=12Q" .